A freight train covered 120 miles at a certain speed. Had the train been able to travel 10 mph faster, the trip would have been 2 hours shorter. How fast did the train go?

Answer by stanbon(75887)

A freight train covered 120 miles at a certain speed. Had the train been able to travel 10 mph faster, the trip would have been 2 hours shorter. How fast did the train go?
-----
slower DATA:
dist = 120 miles ; rate = x mph ; time = 120/x hrs
----------------
faster DATA:
dist = 120 miles ; rate = x+10 mph ; time = 120/(x+10) hrs
-------------------
Equation:
longer time - shorter time = 2hr
120/x - 120/(x+10) = 2 hr
120x + 120*10 - 120x = 2x(x+10)
2x^2 + 20x - 120*10 = 0
------
x^2 + 10x - 600 = 0
x = 20 mph (slower rate)
x+10 = 30 (faster rate)
-----
Cheers,
Stan H.
-------